# A General Model for Linearly Polarized Optical Vector Beams

### v1

@@ I Introduction
$\mathcal{L}$ $=\left[\dot{V}({\bf x},t)\dot{V}^{*}({\bf x},t)-\nabla V({\bf x},t)\cdot\nabla V^{*}({\bf x},t)\right]$ (1)
where the overdot denotes time derivative, ()$\nabla(\cdot)\equiv\{\partial_{x}(\cdot),\partial_{y}(\cdot),\partial_{z}(\cdot)\}$, and $(\cdot)^{*}$ takes the complex conjugate. The authors went on to show that in a homogeneous, isotropic medium, $V({\bf x},t)$ obeyed the familiar scalar wave equation,
@@ II.1 Solutions to the Paraxial Wave Equation
$\qquad\qquad\rho^{1/2}(\vec{x},z)e^{-i\phi(\vec{x},z)}\left\{\cos{\gamma(\vec{x},z)},\sin{\gamma(\vec{x},z)}\right\}$ (5)
@@ II.2 Stationary Action Solutions for Vector Beams
$\mathcal{L}$ $=-\epsilon_{0}k_{0}\rho\partial_{z}\phi-\frac{\epsilon_{0}}{2}\rho|\nabla_{X}\phi)|^{2}-\frac{\epsilon_{0}|\nabla_{X}\rho|^{2}}{8\rho},$ (10)
where $\epsilon_{0}$ is the vacuum permittivity. Note that in this monochromatic model, time has been effectively removed from the problem as the amplitudes depend only on the three spatial coordinates, hence the Lagrangian density is in units $J/m^{3}$. The overdot for the amplitudes in this case represent derivatives w.r.t. $z$.
@@
$k_{0}\frac{D\gamma(\vec{x},z)}{Dz}$ $+k_{0}\partial_{z}\phi(\vec{x},z)+\frac{1}{2}|\nabla_{X}\phi(\vec{x},z)|^{2}+\frac{1}{2}|\nabla_{X}\gamma(\vec{x},z)|^{2}$
$=\frac{1}{2}\frac{\nabla_{X}\rho^{1/2}(\vec{x},z)}{\rho^{1/2}(\vec{x},z)}$ (13)
@@
in the complex paraxial Lagrangian (10) yields the same model (6) as we obtained by substituting the full vector electric field into (3). The resulting system of equations is appropriate for describing the evolution of magnitude, phase, and linear polarization angle in paraxial vector beams.
@@ II.3 Consideration of Momentum Density
$=\rho^{1/2}e^{-i(\phi(\vec{x},z)+k_{0}z-\omega t)}\{\cos(\gamma(\vec{x},z),\sin(\gamma(\vec{x},z)\}$ (17)
@@
${\bf P}_{V}$ $=-\frac{1}{2ck_{0}^{2}}\left[\dot{V}^{*}({\bf x},t)\nabla V({\bf x},t)+\dot{V}({\bf x},t)\nabla V^{*}({\bf x},t)\right].$ (21)
@@ III Proposed Scalar Potential
the Lagrangian energy density in $J/m^{4}$ (note the time operator is normalized by $c$ and hence is converted to a distance) can then be written in several compact forms, all equivalent to (1):
@@
The model (27) is the generalization of (6) where the spatial dependence on the propagation direction $z$ has been replaced with dependence in time while the transverse gradient operators become full, three-dimensional gradients. The “velocity” $\vec{v}$ and polarization gradient terms now have the units $m/s$ (as opposed to being dimensionless quantities in Eqns. 6) and the presence of a medium with $n({\bf x})\neq 1$ is also included.

### v3

@@ I Introduction
$\mathcal{L}$ $=\frac{\epsilon_{0}}{2}\left[\dot{V}({\bf x},t)\dot{V}^{*}({\bf x},t)-\nabla V({\bf x},t)\cdot\nabla V^{*}({\bf x},t)\right]$ (1)
where the overdot denotes time derivative, ()$\nabla(\cdot)\equiv\{\partial_{x}(\cdot),\partial_{y}(\cdot),\partial_{z}(\cdot)\}$, the operator $(\cdot)^{*}$ takes the complex conjugate, and $\epsilon_{0}$ is the vacuum permittivity. The authors went on to show that in a homogeneous, isotropic medium, $V({\bf x},t)$ obeyed the familiar scalar wave equation,
@@ II.1 Solutions to the Paraxial Wave Equation
$\qquad\rho^{1/2}(\vec{x},z)e^{-i\phi(\vec{x},z)}\left\{\cos{\gamma(\vec{x},z)},\sin{\gamma(\vec{x},z)}\right\}$ (5)
@@ II.2 Stationary Action Solutions for Vector Beams
$\mathcal{L}$ $=-\epsilon_{0}k_{0}\rho\partial_{z}\phi-\frac{\epsilon_{0}}{2}\rho|\nabla_{X}\phi|^{2}-\frac{\epsilon_{0}|\nabla_{X}\rho|^{2}}{8\rho},$ (10)
Note that in this monochromatic model, time has been effectively removed from the problem as the amplitudes depend only on the three spatial coordinates, hence the Lagrangian density is in units $J/m^{3}$. The overdot for the amplitudes in this case represent derivatives w.r.t. $z$.
@@
$k_{0}\frac{D\gamma(\vec{x},z)}{Dz}$ $+k_{0}\partial_{z}\phi(\vec{x},z)+\frac{1}{2}|\nabla_{X}\phi(\vec{x},z)|^{2}$
$+\frac{1}{2}|\nabla_{X}\gamma(\vec{x},z)|^{2}=\frac{1}{2}\frac{\nabla_{X}\rho^{1/2}(\vec{x},z)}{\rho^{1/2}(\vec{x},z)}$ (13)
@@
in place of the scalar electric field in the complex paraxial Lagrangian (10) yields the same model (6) as we obtained by substituting the full vector electric field into (3). The resulting system of equations is appropriate for describing the evolution of magnitude, phase, and linear polarization angle in paraxial vector beams.
@@ II.3 Consideration of Momentum Density
$~{}~{}~{}=\rho^{1/2}e^{-i(\phi(\vec{x},z)+k_{0}z-\omega t)}\{\cos(\gamma(\vec{x},z),\sin(\gamma(\vec{x},z)\}$ (17)
@@
${\bf P}_{V}$ $=-\frac{1}{2c}\left[\dot{V}^{*}({\bf x},t)\nabla V({\bf x},t)+\dot{V}({\bf x},t)\nabla V^{*}({\bf x},t)\right].$ (21)
@@ III Proposed Scalar Potential
the Lagrangian energy density in $J/m^{3}$ (note the time operator is normalized by $c$ and hence is converted to a distance) can then be written in several compact forms, all equivalent to (1):
@@
The model (27) is the generalization of (6) where the spatial dependence on the propagation direction $z$ has been replaced with dependence in time while the transverse gradient operators become full, three-dimensional gradients. The “velocity” $\vec{v}$ and polarization gradient terms now have the units $m/s$ (as opposed to being dimensionless quantities in Eqns. 6).